# Revisiting Entropy in Reinforcement Learning for Large Reasoning Models

### v1

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) has emerged as a predominant approach for enhancing the reasoning capabilities of large language models (LLMs). However, the entropy of LLMs usually collapses during RLVR training, causing premature convergence to suboptimal local minima and hinder further performance improvement. Although various approaches have been proposed to mitigate entropy collapse, a comprehensive study of entropy in RLVR remains lacking. To address this gap, we conduct extensive experiments to investigate the entropy dynamics of LLMs trained with RLVR and analyze how model entropy correlates with response diversity, calibration, and performance across various benchmarks. Our findings reveal that the number of off-policy updates, the diversity of training data, and the clipping thresholds in the optimization objective are critical factors influencing the entropy of LLMs trained with RLVR. Moreover, we theoretically and empirically demonstrate that tokens with positive advantages are the primary contributors to entropy collapse, and that model entropy can be effectively regulated by adjusting the relative loss weights of tokens with positive and negative advantages during training.
@@ 1 Introduction
Pioneered by OpenAI o1 (DBLP:journals/corr/abs-2412-16720), DeepSeek-R1 (DBLP:journals/corr/abs-2501-12948), and Kimi k1.5 (DBLP:journals/corr/abs-2501-12948), reinforcement learning with verifiable rewards (RLVR) has been widely employed to extend the boundaries of reasoning capacity in large language models (LLMs). LLMs trained with RLVR have demonstrated remarkable performance on tasks requiring sophisticated reasoning processes, such as mathematics and coding DBLP:journals/corr/abs-2502-03387; DBLP:journals/corr/abs-2503-20783; deepcoder2025; DBLP:journals/corr/abs-2505-22312.
@@
However, although RLVR enhances the reasoning ability of LLMs, a growing body of research has shown that it can also drive LLMs toward entropy collapse, wherein the entropy of the model decreases substantially during training, ultimately reaching a markedly low level (DBLP:journals/corr/abs-2503-14476; DBLP:journals/corr/abs-2508-11016). Entropy collapse indicates that the probability mass over the model’s vocabulary becomes concentrated on a limited subset of tokens. This phenomenon further implies that LLMs increasingly prioritize exploitation over exploration. As a result, they fail to effectively explore novel reasoning paths during training, thereby potentially causing premature convergence to a local optimum.
@@
To address entropy collapse in LLMs, numerous methods have been proposed. shen2025entropycontrolllmrlalgorithms and DBLP:journals/corr/abs-2505-22312 incorporate an entropy maximization term into the RLVR objective. DAPO (DBLP:journals/corr/abs-2503-14476) raises the upper clipping bound of the importance ratio to avoid clipping low-probability tokens. DBLP:journals/corr/abs-2505-22617 restrict parameter updates for tokens with high covariance between log-probability and advantage, alongside other approaches (DBLP:journals/corr/abs-2506-01939; DBLP:journals/corr/abs-2508-10751; DBLP:journals/corr/abs-2508-02260; DBLP:journals/corr/abs-2508-11016; DBLP:journals/corr/abs-2506-01347). Despite these advances, systematic investigations of entropy in RLVR remain scarce. Specifically, three critical questions are still underexplored: (1) How does the entropy of LLMs trained with RLVR correlate with their performance? (2) What factors govern entropy dynamics, both theoretically and empirically? and (3) How can entropy be effectively regulated to improve the performance of LLMs?
@@
To investigate the above research questions, we conduct extensive experiments on RLVR. The results show that the entropy of LLMs trained with RLVR is strongly correlated with response diversity, as LLMs with lower entropy tend to produce less diverse outputs. During training, both response and prompt entropy consistently decline, with in-domain prompts decreasing more rapidly than out-of-domain ones. However, prompt entropy shows only a weak correlation with the accuracy of corresponding responses. Moreover, we observe that model performance can continue to improve even without a reduction in entropy. Entropy also proves to be an unreliable indicator of model performance across most benchmarks, as the correlations between entropy and performance are task-dependent. Empirically, we find that entropy collapse is associated with model miscalibration, and that more severe entropy collapse corresponds to stronger miscalibration. Beyond these findings, we identify three key factors that influence entropy dynamics: (1) the number of off-policy updates, (2) training data diversity, and (3) the clipping threshold. Surprisingly, LLMs trained on approximately 600 samples achieve performance comparable to those trained on about 17k samples. Finally, we theoretically and empirically demonstrate that tokens with positive advantages are the primary contributors to entropy collapse. Adjusting the relative loss weights assigned to tokens with positive and negative advantages effectively regulates entropy and improves LLM performance on benchmarks. Our contributions can be summarized as follows:
@@
  • We conduct extensive experiments to investigate the entropy dynamics of LLMs trained with RLVR, revealing how entropy interacts with response diversity, calibration, and performance on various benchmarks.
@@
  • We identify three key factors influencing the entropy dynamics of LLMs during RLVR training: (1) off-policy updates, (2) training data diversity, and (3) clipping threshold.
@@
  • We theoretically and empirically demonstrate that entropy collapse in RLVR primarily arises from tokens with positive advantages, and that regulating their relative loss weights provides an effective means to control entropy and improve model performance.
@@ 2 Related Work
Entropy has long served as a regularization mechanism to encourage exploration in reinforcement learning (DBLP:conf/aaai/ZiebartMBD08; DBLP:phd/us/Ziebart18). In the era of LLMs, several studies have similarly incorporated entropy maximization objectives into RLVR to promote exploration and prevent premature convergence (shen2025entropycontrolllmrlalgorithms; DBLP:journals/corr/abs-2505-22312). Aiming to mitigate entropy collapse, Clip-Higher (DBLP:journals/corr/abs-2503-14476) raises the upper clipping bound of the importance ratio to prevent low-probability tokens with positive advantages from being clipped. zhihu2025entropy and DBLP:journals/corr/abs-2505-22617 provide theoretical insights into entropy dynamics, showing that tokens with strong positive covariance between their probabilities and corresponding advantages primarily drive entropy collapse. Building on these insights, Clip-Cov and KL-Cov (DBLP:journals/corr/abs-2505-22617) limit updates on tokens exhibiting such covariance. Similarly, CE-GPPO (su2025cegppocoordinatingentropygradientpreserving) mitigates entropy collapse by preserving the gradients of clipped tokens through a stop-gradient operation. Furthermore, DBLP:journals/corr/abs-2506-01939 train LLMs using only high-entropy tokens to enhance model performance, while DBLP:journals/corr/abs-2506-14758 incorporate entropy terms into the advantage to encourage exploration and improve reasoning capabilities. Numerous other studies have also explored entropy-based mechanisms to further enhance the performance of LLMs trained with RLVR (DBLP:journals/corr/abs-2508-11016; wang2025harnessinguncertaintyentropymodulatedpolicy; liu2025uniformheterogeneoustailoringpolicy).

### v2

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) has emerged as a prominent paradigm for enhancing the reasoning capabilities of large language models (LLMs). However, the entropy of LLMs usually collapses during RLVR training, leading to premature convergence to suboptimal local minima and hindering further performance improvement. Although various approaches have been proposed to mitigate entropy collapse, a comprehensive study of entropy in RLVR remains lacking. To bridge this gap, we conduct extensive experiments to investigate the entropy dynamics of LLMs trained with RLVR and analyze how model entropy correlates with response diversity, calibration, and performance across various benchmarks. Our results identify three key factors that influence entropy: the clipping thresholds in the optimization objective, the number of off-policy updates, and the diversity of the training data. Furthermore, through both theoretical analysis and empirical validation, we demonstrate that tokens with positive advantages are the primary drivers of entropy collapse. Motivated by this insight, we propose Positive-Advantage Reweighting, a simple yet effective approach that regulates model entropy by adjusting the loss weights assigned to tokens with positive advantages during RLVR training, while maintaining competitive performance.
@@ 1 Introduction
Pioneered by OpenAI o1 (Jaech et al., 2024), DeepSeek-R1 (DeepSeek-AI et al., 2025), and Kimi k1.5 (Team et al., 2025), reinforcement learning with verifiable rewards (RLVR) has been widely employed to extend the boundaries of reasoning capacity in large language models (LLMs). LLMs trained with RLVR have demonstrated remarkable performance on tasks requiring sophisticated reasoning processes, such as mathematics and coding Ye et al. (2025); Liu et al. (2025b); Luo et al. (2025); He et al. (2025).
@@
However, although RLVR enhances the reasoning ability of LLMs, a growing body of research has shown that it can also drive LLMs toward entropy collapse, wherein the entropy of the model decreases substantially during training, ultimately reaching a markedly low level (Yu et al., 2025; Li et al., 2025a). Entropy collapse indicates that the probability mass over the model’s vocabulary becomes concentrated on a limited subset of tokens. This phenomenon further implies that LLMs increasingly prioritize exploitation over exploration. As a result, they fail to effectively explore novel reasoning paths during training, thereby potentially causing premature convergence to a local optimum.
@@
To address entropy collapse in LLMs, numerous methods have been proposed. Shen (2025) and He et al. (2025) incorporate an entropy maximization term into the RLVR objective. DAPO (Yu et al., 2025) raises the upper clipping bound of the importance ratio to avoid clipping low-probability tokens. Cui et al. (2025) restrict parameter updates for tokens with high covariance between log-probability and advantage, alongside other approaches (Wang et al., 2025b; Chen et al., 2025; Deng et al., 2025; Li et al., 2025a; Zhu et al., 2025). Despite these advances, systematic investigations of entropy in RLVR remain scarce. Specifically, three critical questions are still underexplored: (1) How does the entropy of LLMs trained with RLVR correlate with their performance? (2) What factors govern entropy dynamics, both theoretically and empirically? and (3) How can entropy be effectively regulated to improve the performance of LLMs?
@@
To investigate the above research questions, we conduct extensive experiments on RLVR. We find that the entropy of LLMs trained with RLVR is strongly correlated with response diversity, as LLMs with lower entropy tend to produce less diverse outputs. During training, both response entropy and prompt entropy decrease, with in-domain prompt entropy declining more rapidly than that of out-of-domain prompts. Meanwhile, prompt entropy exhibits only a weak correlation with the accuracy of the corresponding responses. Notably, we observe that model performance can continue to improve without sacrificing entropy. In addition, entropy does not serve as a reliable proxy for model performance across most benchmarks, as the observed correlations between entropy and performance are highly task-dependent. We further observe that entropy collapse is closely associated with model miscalibration, and that more severe entropy collapse corresponds to stronger miscalibration. Beyond these empirical findings, we identify three key factors influencing entropy dynamics: (1) the clipping threshold, (2) the number of off-policy updates, and (3) training data diversity. Remarkably, an LLM trained on approximately 600 samples can achieve performance comparable to one trained on around 17k samples. Finally, through both theoretical analysis and empirical validation, we demonstrate that tokens with positive advantages are the primary drivers of entropy collapse. Motivated by this insight, we propose Positive-Advantage Reweighting, which adjusts the loss weights of tokens with positive advantages to regulate model entropy while improving performance. Our contributions can be summarized as follows:
@@
  • We conduct extensive experiments to investigate the dynamics of prompt entropy and response entropy in LLMs trained with RLVR, revealing how entropy correlates with response diversity, model calibration, and performance.
@@
  • We identify three key factors influencing the entropy dynamics of LLMs during RLVR training: (1) clipping threshold, (2) off-policy updates, and (3) training data diversity.
@@
  • We theoretically and empirically demonstrate that entropy collapse in RLVR primarily arises from positive-advantage tokens, and propose Positive-Advantage Reweighting to control LLM entropy and improve performance by dynamically reweighting the losses of such tokens.
@@ 2 Related Work
Entropy has long served as a regularization mechanism to encourage exploration in reinforcement learning (Ziebart et al., 2008; Ziebart, 2010). In the era of LLMs, several studies have similarly incorporated entropy maximization objectives into RLVR to promote exploration and prevent premature convergence (Shen, 2025; He et al., 2025). Aiming to mitigate entropy collapse, Clip-Higher (Yu et al., 2025) raises the upper clipping bound of the importance ratio to prevent low-probability tokens with positive advantages from being clipped. Liu (2025) and Cui et al. (2025) provide theoretical insights into entropy dynamics, showing that tokens with strong positive covariance between their probabilities and corresponding advantages primarily drive entropy collapse. Building on these insights, Clip-Cov and KL-Cov (Cui et al., 2025) limit updates on tokens exhibiting such covariance. Similarly, CE-GPPO (Su et al., 2025) mitigates entropy collapse by preserving the gradients of clipped tokens through a stop-gradient operation. Furthermore, Wang et al. (2025b) train LLMs using only high-entropy tokens to enhance model performance, while Cheng et al. (2025) incorporate entropy terms into the advantage to encourage exploration and improve reasoning capabilities. Numerous other studies have also explored entropy-based mechanisms to further enhance the performance of LLMs trained with RLVR (Li et al., 2025a; Wang et al., 2025a; Liu et al., 2025a).